# Lost in the FoG: Pitfalls of Models for Large-Scale Hydrogen Distributions

### v4

@@ 4.3 Redshift-space distortions
Figure 5: The 2D auto power spectra of all matter (top center), ${\rm H}$i (top right), and blue (bottom left), red (bottom center), and all-galaxies (bottom right) at $z=1$. Black dotted lines are concentric circles to help visualize the FoG effect, which vertically warps the isopower contours. This warping is strongest in red galaxies as compared to the other galaxy types, agreeing with observations (Madgwick et al., 2003). We find that ${\rm H}$i appears to have stronger FoG suppression than matter, as found in VN18, although their relative strengths depend on scale (see text). Power spectra are slightly smoothed by a Gaussian filter ($\sigma=0.5$) to see the contour lines more clearly, but the smoothed values are not used for the determination of the pairwise velocity dispersion and do not otherwise impact our conclusions.
@@
Note. — Pairwise velocity dispersion (PVD) values retrieved from fits of the FoG damping term to the 2D power spectrum presented in Fig. 5. We fit down to $k\leq 1$ h cMpc-1, which limits shot noise contributions. The $\chi^{2}$ values are calculated over the same scale regime. The galaxy populations generally have large residuals — we tested different $k$ limits and found no significant improvement in $\chi^{2}$, although the PVD values change by $\lesssim 0.1$. We attribute the poor fits to absent nonlinearities and scale-dependent effects in the FoG term, which has been found in other work (Ando et al., 2019).
@@
The FoG effect suppresses clustering at small scales, elongating the contour lines in the top left corners of each panel in Fig. 5. Amongst the galactic populations in the bottom row, the contour lines in the red galaxy panel are warped the furthest vertically, agreeing qualitatively with observations (Madgwick et al., 2003). The red galaxies’ large FoG suppression originates from their tendency to occupy massive halos with large velocity dispersions (Zehavi et al., 2011). Conversely, blue galaxies occupy smaller halos, resulting in a weaker FoG effect. The all-galaxies’ (bottom right panel) FoG strength falls between blue and red galaxies’, as expected for a combined sample. Of the two remaining populations in the top row of Fig. 5, the FoG damping in ${\rm H}$i appears to be stronger than that of matter, as seen from the more vertical contour lines in the panel’s top right corner. This agrees with conclusions from VN18, but the following discussion about Table 3 will demonstrate that this finding does not hold quantitatively.
We measure the strength of the FoG effect with the “pairwise velocity dispersion” (PVD, see Section 3). This parameter is difficult to measure directly (e.g., Loveday et al., 2018), particularly in intensity maps which do not observe point sources. Instead, the PVD is usually left as a free parameter to marginalize over (e.g., Gil-Marín et al., 2020; de Mattia et al., 2021). Following VN18, we extract the PVD (denoted $\sigma_{p}$) by fitting the following equation to the 2D power spectra of the population $i$ on scales $k\leq 1$ h cMpc-1:
@@
Some of the tension between Fig. 5 and Table 3 can be credited to the poor fits, which we quantify with the $\chi^{2}$ values in Table 3. ${\rm H}$i and matter generally maintain small $\chi^{2}$ values, but the galaxy populations all contain large values. To ensure that the poor fits originate from the model rather than our methodology, we tested the sensitivity of the $\chi^{2}$ to the scales we included in our fits ($k\leq 1$ h cMpc-1). All tests produced larger $\chi^{2}$, although we found that the scale regime could change the PVD values by $\sim$ 0.1. PVD is a scale-dependent quantity (see Section 3.5), so the sensitivity to scale is expected. These tests suggest the poor fits stem from representing RSDs with Equation 11 for blue and red galaxy distributions, which agrees with results from other works (e.g., Hikage & Yamamoto, 2015; Orsi & Angulo, 2018).
@@
[…]
to the actual redshift-space power spectrum for each population in Fig. 6 (additional redshifts provided in the online figures). Generally, we find that the integrated 2D power spectra overestimate the actual redshift-space power spectra for each population, and particularly for galaxies. ${\rm H}$i is the earliest tracer to fall below unity at $k\sim 0.4$ h cMpc-1, although all of the galaxy tracers follow suit around $k\sim 0.9$ h cMpc-1. We attribute this feature in the galaxy ratios to the shot noise floor that they reach around that scale, visible in their redshift-space power spectra (online figures). We limit the impact of shot noise by restricting our fits to $k\leq 1$ h cMpc-1, but as previously mentioned, the fits do not improve with more or less conservative $k$ limits.
In summary, we provide the 2D power spectra for matter and each tracer in Fig. 5, and fit PVD values to them using the phenomenological model from Equation 15. However, we find that these fits misrepresent the 1D power spectrum when integrated, which may arise from the particular functional form in Equation 11 as we speculate in Section 6.2. For now, we will broadly refer to the differences shown in Fig. 6 as “RSD errors”.
@@ 5 Testing the model predictions
In Section 4, we discussed each ingredient of the general model in Equations 12-13 for ${\rm H}$i auto and cross-power spectra. The goal of this section is to understand the aggregate effect of each ingredient on the model’s performance by comparing their predictions to the simulated power spectra, seeking to answer the question “Given the ${\rm H}$i distribution in IllustrisTNG, how accurately could these models extract cosmological parameters from a perfectly measured power spectrum?” We know that any deviations between the models and simulated power spectra must arise from the combination of these ingredients failing to capture tracer clustering behavior, since each ingredient is extracted directly from the simulation. Moreover, these ingredients will be realistically further degraded by systematics and noise in observations, which are not present in our analyses. As such, we emphasize that our tests represent the best-case scenario and stated errors should be taken as lower limits. Qualitatively, the following results should be robust to cosmic variance although second-order effects could change some details (e.g., Li et al., 2014).
5.1 Hi auto power spectra
@@
The ${\rm H}$i auto power spectrum in IllustrisTNG agrees closely with recent observations of the $z\approx 0.5$ ${\rm H}$i auto power spectrum (P23; Osinga et al., 2024), presenting a great opportunity to leverage IllustrisTNG to test common models of large-scale ${\rm H}$i distributions. In Fig. 7, we compare four models of varying complexity by showing their ratios over the actual $z=0.5$ redshift-space ${\rm H}$i auto power spectrum from IllustrisTNG, with darker colors corresponding to more complex models (plots for additional redshifts are provided in Appendix C). The details of each prescription are described in the legend, with the background color of each row matching the corresponding power spectrum. The first two columns of the legend indicate whether or not the model assumes a constant bias or neglects the FoG effect. The third column details the equivalent equation, and the fourth column includes a shorthand that we use in the text.
@@
On the largest scales, we expect small differences between the various models since the assumptions made by the simpler prescriptions should be appropriate (see Section 3.5). While all models converge on the largest scales, interestingly they do not converge to the actual redshift-space ${\rm H}$i auto power spectrum, overestimating it by a factor of 1.08. Noise emerging from a small number of large-scale could contribute but is unlikely to account for this difference entirely. We instead attribute the large-scale offset to the Kaiser term since all models err by the same amount regardless of bias or FoG treatment. The Kaiser term neglects nonlinearities in the matter velocity fields, which are known to emerge at $k\sim 0.05$ h cMpc-1 (Reid & White, 2011; Jennings et al., 2011; Howlett et al., 2017). Despite the departures on the largest scales, all models remain within 10% of the actual power spectrum at $k\lesssim 0.15$ h cMpc-1.
@@
The two simplest prescriptions cross the $10\%$ threshold at similar scales, whether a constant (cB) or scale-dependent (sB) bias is assumed. The small difference in performance between cB and sB, despite the impractically precise description of the ${\rm H}$i bias in sB, suggests that including non-constant bias terms in ${\rm H}$i auto power spectra models would negligibly improve their performance. The errors in these models worsen to $>25\%$ on the scales relevant to a recent observation of the ${\rm H}$i auto power spectrum, P23222We note that P23 remove $k_{\parallel}\leq 0.3k_{\perp}$ modes from their analysis due to instrumentation effects, which should change model performance. As such, Fig. 7 should be interpreted as the model errors from a similar survey without this limitation. (green shaded area). Most of the error from the cB and sB models arises from neglecting the FoG term, as shown by the significant improvement in cB+FoG. Despite neglecting non-constant terms in the ${\rm H}$i bias, the cB+FoG error on observation scales remains at $\lesssim 15\%$ until $k\sim 1.5$ h cMpc-1.
@@
In summary, Fig. 7 shows that neglecting the FoG term, even on scales $k\sim 0.1$ h cMpc-1, causes inaccuracies that dominate other potential error sources, like neglecting non-constant terms in the ${\rm H}$i bias. We also demonstrate that competing error sources can reduce the net error of a simple model, to the point it appears more “accurate” than a more complex model. However, even the full prescription incurs an error of $>10\%$ on observationally relevant scales, which would significantly skew the cosmological constraints inferred from ${\rm H}$i auto power spectra.
@@ 5.2 Hi-galaxy cross-power spectra
Figure 8: Accuracy of models of ${\rm H}$i $\times$ Blue (left), ${\rm H}$i $\times$ Red (center), and ${\rm H}$i $\times$ Galaxy (right) at $z=1$, compared to IllustrisTNG. Darker colors correspond to more complex models (see legend). Green areas show the typical scales of observations. The simplest prescription represents the model used to interpret recent observations, which errs by $\sim 25\%$ even to the largest scales probed by TNG300. Better performing models stay near unity to smaller scales. As expected, more complex models capture more of the behavior of their respective cross-power spectra. However, even the full prescription for ${\rm H}$i $\times$ Red errs by $>50\%$ at $k\sim 1$ h cMpc-1.
Cross-correlations between ${\rm H}$i and galaxy surveys reduce the systematics and noise present in the ${\rm H}$i auto power spectrum, at the cost of additional complexity (Equations 12-13). In this section, we investigate model errors of ${\rm H}$i-galaxy cross-power spectra in Fig. 8 at $z=1$, coinciding with most observations (Chang et al., 2010; Masui et al., 2013; Wolz et al., 2022; CHIME Collaboration et al., 2023). Additional redshifts are provided in Appendix C.
[…]
@@ 5.3 Effect on inferred $\Omega_{\mathrm{HI}}$
Note. — The $\Omega_{\mathrm{HI}}$ constraints from ${\rm H}$i-galaxy cross-power spectra reported in W22, C22, CHIME23, and C24, and ${\rm H}$i auto power spectra from P23. The galaxy survey indicates the cross-correlated galaxy population: WiggleZ (Blake et al., 2011) and ELG (Raichoor et al., 2020) and LRG (Ross et al., 2020) samples from the eBOSS survey. $z_{\mathrm{eff}}$ and $k_{\mathrm{eff}}$ represent the effective redshift and wavenumber (in h cMpc-1) the measurement takes place. The models listed are the closest analog to the model adopted in each work. $10^{3}\mathcal{A}_{\mathrm{HI}}$ is the quantity directly measured by W22, C22, and C24. CHIME23 reported $\mathcal{A}_{\mathrm{HI}}=10^{3}\Omega_{\mathrm{HI}}(b_{\mathrm{HI}}+\langle f\mu^{2}\rangle)$, which we adjust to match Equation 20. These works then assume ${\rm H}$i bias ($b_{\mathrm{HI}}$) and ${\rm H}$i-galaxy correlation coefficient ($r_{\mathrm{HI-Gal}}$) values to roughly estimate $10^{3}\Omega_{\mathrm{HI}}$ from $\mathcal{A}_{\mathrm{HI}}$, sometimes incorporating uncertainties from the assumed values into their $\Omega_{\mathrm{HI}}$ constraints. C24 is an extension of results from C22, so we employ the same assumed values although C24 does not directly make such assumptions. Conversely, P23 fit several parameters to their measured ${\rm H}$i auto power spectra and thus do not assume any values, but they do report $\Omega_{\mathrm{HI}}$ values with and without priors in their fits. We adopt their $\Omega_{\mathrm{HI}}$ values from those with priors since they agree more closely with other constraints, but we caution that the uncertainties in $\Omega_{\mathrm{HI}}$ are likely underestimated.
5.3.1 Accuracy of $\Omega_{\mathrm{HI}}$ constraints from power spectra
Measurements of $\Omega_{\mathrm{HI}}$ probe the state of star-formation and gas in galaxies over cosmic time and inform galaxy formation models (Wyithe, 2008; Diemer et al., 2019; Davé et al., 2020). 21cm intensity mapping experiments can determine $\Omega_{\mathrm{HI}}$ via the amplitude of ${\rm H}$i power spectra since they directly measure brightness temperature fluctuations of ${\rm H}$i, which rescale the density fluctuations by the mean ${\rm H}$i brightness temperature (Battye et al., 2013):
@@ 5.3.1 Accuracy of $\Omega_{\mathrm{HI}}$ constraints from power spectra
for the auto and cross-power spectra, respectively. We vary the minimum scale included in the fit to examine the small-scale trade-off between the larger errors and smaller uncertainties, expressing the inferred $\Omega_{\mathrm{HI}}$ values as a function of $k_{\mathrm{eff}}$. We follow previous literature to estimate the power spectra uncertainties used in the fits (Furlanetto & Lidz, 2007; Lidz et al., 2009; Park et al., 2014), removing any instrumental contributions. We compare the value determined in the fit, $\Omega_{\rm{HI,fit}}$, to $\Omega_{\mathrm{HI,TNG}}$, providing the results for each model of the ${\rm H}$i auto and cross-power spectrum in Fig. 9. We exclude error bars from uncertainties in the fitted values in Fig .9 for clarity (error bars included in online figures), but we note that all uncertainties are negligible except for the value at the smallest $k_{\mathrm{eff}}$.
Ideally, the $\Omega_{\mathrm{HI,fit}}/\Omega_{\mathrm{HI,TNG}}$ ratios of all models should approach unity on large scales, given that the largest scales probed by TNG300 should be linear (Smith et al., 2003). The models in Fig. 9 do converge on large scales, with one exception: mK+cB, which at $k_{\mathrm{eff}}\approx 0.05$ h cMpc-1 only reaches a ratio of about 0.8 and 0.9 for $z=1$ and $z=0.5$, respectively. These errors would significantly skew the recent constraints of $\Omega_{\mathrm{HI}}$ that resorted to this model (see Section 5.3.2). Interestingly, the remaining models also do not converge quite to unity, instead converging to $\approx 0.97$ and $\approx 0.94$ for the auto and cross-power spectra, respectively. These large-scale differences may imply that there is an absent large-scale effect in the models, although we reserve further analysis to future work with a larger-volume simulation.

### v5

@@ 4.3 Redshift-space distortions
Note. — Pairwise velocity dispersion (PVD) values retrieved from fits of the FoG damping term to the 2D power spectrum presented in Fig. 5. We fit down to $k\leq 1$ h cMpc-1, which limits shot noise contributions. The $\chi^{2}$ values are calculated over the same scale regime. The galaxy populations generally have large residuals — we tested different $k$ limits and found no significant improvement in $\chi^{2}$, although the PVD values change by $\lesssim 0.1$. We attribute the poor fits to absent nonlinearities and scale-dependent effects in the FoG term, which has been found in other work (R. Ando et al., 2019).
Figure 5: The 2D auto power spectra of all matter (top center), $\rm H$}i (top right), and blue (bottom left), red (bottom center), and all-galaxies (bottom right) at $z=1$. Black dotted lines are concentric circles to help visualize the FoG effect, which vertically warps the isopower contours. This warping is strongest in red galaxies as compared to the other galaxy types, agreeing with observations (D. S. Madgwick et al., 2003). We find that $\rm H$}i appears to have stronger FoG suppression than matter, as found in VN18, although their relative strengths depend on scale (see text). Power spectra are slightly smoothed by a Gaussian filter ($\sigma=0.5$) to see the contour lines more clearly, but the smoothed values are not used for the determination of the pairwise velocity dispersion and do not otherwise impact our conclusions.
@@
The FoG effect suppresses clustering at small scales, elongating the contour lines in the top left corners of each panel in Fig. 5. Amongst the galactic populations in the bottom row, the contour lines in the red galaxy panel are warped the furthest vertically, agreeing qualitatively with observations (D. S. Madgwick et al., 2003). The red galaxies’ large FoG suppression originates from their tendency to occupy massive halos with large velocity dispersions (I. Zehavi et al., 2011). Conversely, blue galaxies occupy smaller halos, resulting in a weaker FoG effect. The all-galaxies’ (bottom right panel) FoG strength falls between blue and red galaxies’, as expected for a combined sample. Of the two remaining populations in the top row of Fig. 5, the FoG damping in $\rm H$}i appears to be stronger than that of matter, as seen from the more vertical contour lines in the panel’s top right corner. This agrees with conclusions from VN18, but the following discussion about Table 3 will demonstrate that this finding does not hold quantitatively.
We measure the strength of the FoG effect with the “pairwise velocity dispersion” (PVD, see Section 3). This parameter is difficult to measure directly (e.g., J. Loveday et al., 2018), particularly in intensity maps which do not observe point sources. Instead, the PVD is usually left as a free parameter to marginalize over (e.g., H. Gil-Marín et al., 2020; A. de Mattia et al., 2021). Following VN18, we extract the PVD (denoted $\sigma_{p}$) by fitting the following equation to the 2D power spectra of the population $i$ on scales $k\leq 1$ h cMpc-1:
@@
Some of the tension between Fig. 5 and Table 3 can be credited to the poor fits, which we quantify with the $\chi^{2}$ values in Table 3. $\rm H$}i and matter generally maintain small $\chi^{2}$ values, but the galaxy populations all contain large values. To ensure that the poor fits originate from the model rather than our methodology, we tested the sensitivity of the $\chi^{2}$ to the scales we included in our fits ($k\leq 1$ h cMpc-1). All tests produced larger $\chi^{2}$, although we found that the scale regime could change the PVD values by $\sim$ 0.1. PVD is a scale-dependent quantity (see Section 3.5), so the sensitivity to scale is expected. These tests suggest the poor fits stem from representing RSDs with Equation 11 for blue and red galaxy distributions, which agrees with results from other works (e.g., C. Hikage & K. Yamamoto, 2015; Á. A. Orsi & R. E. Angulo, 2018). We also tested a single Lorentzian (following D. Sarkar & S. Bharadwaj, 2018) to represent the FoG term for the $\rm H$}i auto power spectrum (as compared to the Lorentzian squared from Equation 12), but found no improvement in performance, agreeing with conclusions from VN18.
@@
$P_{\mathrm{1D},i}(k)=\frac{1}{2}\int_{-1}^{1}P_{i}^{S}(k,\mu)d\mu\,$, (16)
to the actual redshift-space power spectrum for each population in Fig. 6 (additional redshifts provided in the online figures). Generally, we find that the integrated 2D power spectra overestimate the actual redshift-space power spectra for each population, and particularly for galaxies. The $\rm H$}i ratio falls below unity at $k\sim 0.4$ h cMpc-1, although all of the galaxy tracers follow suit at smaller scales ($k\sim 0.9$ h cMpc-1). We attribute this feature in the galaxy ratios to the shot noise floor that they reach around that scale, visible in their redshift-space power spectra (online figures). We limit the impact of shot noise by restricting our fits to $k\leq 1$ h cMpc-1, but as previously mentioned, the fits do not improve with more or less conservative $k$ limits.
In summary, we provide the 2D power spectra for matter and each tracer in Fig. 5, and fit PVD values to them using the phenomenological model from Equation 15. However, we find that these fits misrepresent the 1D power spectrum when integrated, which may arise from the particular functional form in Equation 11 as we speculate in Section 6.2. For now, we will broadly refer to the differences shown in Fig. 6 as “RSD errors”. In the following sections, we do not display error bars that arise from uncertainties in the PVD values because they are not statistical in nature (Fig. 6).
@@ 5 Testing the model predictions
In Section 4, we discussed each ingredient of the general model in Equations 12-13 for $\rm H$}i auto and cross-power spectra. The goal of this section is to understand the aggregate effect of each ingredient on the model’s performance by comparing their predictions to the simulated power spectra, seeking to answer the question “Given the $\rm H$}i distribution in IllustrisTNG, how accurately could these models extract cosmological parameters from a perfectly measured power spectrum?” We know that any deviations between the models and simulated power spectra must arise from the combination of these ingredients failing to capture tracer clustering behavior, since each ingredient is extracted directly from the simulation. Moreover, these ingredients will be realistically further degraded by systematics and noise in observations, which are not present in our analyses. As such, we emphasize that our tests represent the best-case scenario and stated errors should be taken as lower limits. Qualitatively, the following results should be robust to cosmic variance although second-order effects could change some details (e.g., Y. Li et al., 2014).
5.1 $\rm H$}i auto power spectra
@@
The $\rm H$}i auto power spectrum in IllustrisTNG agrees closely with recent observations of the $z\approx 0.5$ $\rm H$}i auto power spectrum (P23; C. K. Osinga et al., 2024), presenting a great opportunity to leverage IllustrisTNG to test common models of large-scale $\rm H$}i distributions. In Fig. 7, we compare four models of varying complexity by showing their ratios over the actual $z=0.5$ redshift-space $\rm H$}i auto power spectrum from IllustrisTNG, with darker colors corresponding to more complex models (plots for additional redshifts are provided in Appendix C). The details of each prescription are described in the legend, with the background color of each row matching the corresponding power spectrum. The first two columns of the legend indicate whether or not the model assumes a constant bias or neglects the FoG effect. The third column details the equivalent equation, and the fourth column includes a shorthand that we use in the text.
@@
On the largest scales, we expect small differences between the various models since the assumptions made by the simpler prescriptions should be appropriate (see Section 3.5). While all models converge on the largest scales, interestingly they do not converge to the actual redshift-space $\rm H$}i auto power spectrum, overestimating it by a factor of 1.08. Noise emerging from a small number of large-scale could contribute but is unlikely to account for this difference entirely. We instead attribute the large-scale offset to the Kaiser term since all models err by the same amount regardless of bias or FoG treatment. The Kaiser term neglects nonlinearities in the matter velocity fields, which are known to emerge at $k\sim 0.05$ h cMpc-1 (B. A. Reid & M. White, 2011; E. Jennings et al., 2011; C. Howlett et al., 2017). Despite the departures on the largest scales, all models remain within 10% of the actual power spectrum at $k\lesssim 0.15$ h cMpc-1.
@@
The two simplest prescriptions cross the $10\%$ threshold at similar scales, whether a constant (cB) or scale-dependent (sB) bias is assumed. The small difference in performance between cB and sB, despite the impractically precise description of the $\rm H$}i bias in sB, suggests that including non-constant bias terms in $\rm H$}i auto power spectra models would negligibly improve their performance. The errors in these models worsen to $25\%$ on the scales relevant to a recent observation of the $\rm H$}i auto power spectrum, P23333We note that P23 remove $k_{\parallel}\leq 0.3k_{\perp}$ modes from their analysis due to instrumentation effects, which should change model performance. As such, Fig. 7 should be interpreted as the model errors from a similar survey without this limitation. (green shaded area). Most of the error from the cB and sB models arises from neglecting the FoG term, as shown by the significant improvement in cB+FoG. Despite neglecting non-constant terms in the $\rm H$}i bias, the cB+FoG error on observation scales remains at $\lesssim 15\%$ until $k\sim 1.5$ h cMpc-1.
@@
In summary, Fig. 7 shows that neglecting the FoG term, even on scales $k\sim 0.1$ h cMpc-1, causes inaccuracies that dominate other potential error sources, like neglecting non-constant terms in the $\rm H$}i bias. We also demonstrate that competing error sources can reduce the net error of a simple model, to the point it appears more “accurate” than a more complex model. However, even the full prescription incurs an error of $10\%$ on observationally relevant scales, which would significantly skew the cosmological constraints inferred from $\rm H$}i auto power spectra.
@@ 5.2 Hi-galaxy cross-power spectra
Figure 8: Accuracy of models of $\rm H$}i $\times$ Blue (left), $\rm H$}i $\times$ Red (center), and $\rm H$}i $\times$ Galaxy (right) at $z=1$, compared to IllustrisTNG. Darker colors correspond to more complex models (see legend). Green areas show the typical scales of observations. The simplest prescription represents the model used to interpret recent observations, which errs by $\sim 25\%$ even to the largest scales probed by TNG300. Better performing models stay near unity to smaller scales. As expected, more complex models capture more of the behavior of their respective cross-power spectra. However, even the full prescription for $\rm H$}i $\times$ Red errs by $50\%$ at $k\sim 1$ h cMpc-1.
Cross-correlations between $\rm H$}i and galaxy surveys reduce the systematics and noise present in the $\rm H$}i auto power spectrum, at the cost of additional complexity (Equations 12-13). In this section, we investigate model errors of $\rm H$}i-galaxy cross-power spectra in Fig. 8 at $z=1$, coinciding with most observations (T.-C. Chang et al., 2010; K. W. Masui et al., 2013; L. Wolz et al., 2022; CHIME Collaboration et al., 2023). Additional redshifts are provided in Appendix C.
Although both Fig. 7 and 8 show four models each, the models in the latter differ from the former in two ways. First, recent measurements of the $\rm H$}i-galaxy cross-power spectrum used the matter Kaiser term (Equation 7, $al@wolz_h_2022, cunnington_h_2022, Carucci24; \al@wolz$_h_2022, cunnington_h_2022, Carucci24, ), which we include in Fig. 8. As a result, the naming convention for cross-power spectra models includes a specification for the Kaiser term, whereas auto power spectra only include the type of bias and FoG terms (see legends of Fig. 7 and 8). This additional specification distinguishes the auto and cross-power spectra model shorthands. Second, we remove a cross-power spectra model that uses a scale-dependent bias but neglects the FoG effect (analogous to sB from Section 5.1) since we established in the previous section that neglecting FoG is the dominant source of error.
@@ 5.3 Effect on inferred $\Omega_{\mathrm{HI}}$
5.3.1 Accuracy of $\Omega_{\mathrm{HI}}$ constraints from power spectra
Measurements of $\Omega_{\mathrm{HI}}$ probe the state of star-formation and gas in galaxies over cosmic time and inform galaxy formation models (J. S. B. Wyithe, 2008; B. Diemer et al., 2019; R. Davé et al., 2020). 21cm intensity mapping experiments can determine $\Omega_{\mathrm{HI}}$ via the amplitude of $\rm H$}i power spectra since they directly measure brightness temperature fluctuations of $\rm H$}i, which rescale the density fluctuations by the mean $\rm H$}i brightness temperature (R. A. Battye et al., 2013):
@@ 5.3.1 Accuracy of $\Omega_{\mathrm{HI}}$ constraints from power spectra
Note. — The $\Omega_{\mathrm{HI}}$ constraints from $\rm H$}i-galaxy cross-power spectra reported in W22, C22, CHIME23, and C24, and $\rm H$}i auto power spectra from P23. The galaxy survey indicates the cross-correlated galaxy population: WiggleZ (C. Blake et al., 2011) and ELG (A. Raichoor et al., 2020) and LRG (A. J. Ross et al., 2020) samples from the eBOSS survey. $z_{\mathrm{eff}}$ and $k_{\mathrm{eff}}$ represent the effective redshift and wavenumber (in h cMpc-1) the measurement takes place. The models listed are the closest analog to the model adopted in each work. $10^{3}\mathcal{A}_{\mathrm{HI}}$ is the quantity directly measured by W22, C22, and C24. CHIME23 reported $\mathcal{A}_{\mathrm{HI}}=10^{3}\Omega_{\mathrm{HI}}(b_{\mathrm{HI}}+\langle f\mu^{2}\rangle)$, which we adjust to match Equation 20. These works then assume $\rm H$}i bias ($b_{\mathrm{HI}}$) and $\rm H$}i-galaxy correlation coefficient ($r_{\mathrm{HI-Gal}}$) values to roughly estimate $10^{3}\Omega_{\mathrm{HI}}$ from $\mathcal{A}_{\mathrm{HI}}$, sometimes incorporating uncertainties from the assumed values into their $\Omega_{\mathrm{HI}}$ constraints. C24 is an extension of results from C22, so we employ the same assumed values although C24 does not directly make such assumptions. Conversely, P23 fit several parameters to their measured $\rm H$}i auto power spectra and thus do not assume any values, but they do report $\Omega_{\mathrm{HI}}$ values with and without priors in their fits. We adopt their $\Omega_{\mathrm{HI}}$ values from those with priors since they agree more closely with other constraints, but we caution that the uncertainties in $\Omega_{\mathrm{HI}}$ are likely underestimated.
@@
for the auto and cross-power spectra, respectively. We vary the minimum scale included in the fit to examine the small-scale trade-off between the larger errors and smaller uncertainties, expressing the inferred $\Omega_{\mathrm{HI}}$ values as a function of $k_{\mathrm{eff}}$. We follow previous literature to estimate the power spectra uncertainties used in the fits (S. R. Furlanetto & A. Lidz, 2007; A. Lidz et al., 2009; J. Park et al., 2014), removing any instrumental contributions. We compare the value determined in the fit, $\Omega_{\rm{HI,fit}}$, to $\Omega_{\mathrm{HI,TNG}}$, providing the results for each model of the $\rm H$}i auto and cross-power spectrum in Fig. 9. We exclude error bars from uncertainties in the fitted values in Fig .9 for clarity (error bars included in online figures), but we note that all uncertainties are negligible except for the value at the smallest $k_{\mathrm{eff}}$.
Ideally, the $\Omega_{\mathrm{HI,fit}}/\Omega_{\mathrm{HI,TNG}}$ ratios of all models should approach unity on large scales, given that the largest scales probed by TNG300 should be linear (R. E. Smith et al., 2003). The models in Fig. 9 do converge on large scales, with one exception: mK+cB, which at $k_{\mathrm{eff}}\approx 0.05$ h cMpc-1 only reaches a ratio of about 0.8 and 0.9 for $z=1$ and $z=0.5$, respectively. These errors would significantly skew the recent constraints of $\Omega_{\mathrm{HI}}$ that resorted to this model (see Section 5.3.2). Interestingly, the remaining models also do not converge quite to unity, instead converging to $\approx 0.97$ and $\approx 0.94$ for the auto and cross-power spectra, respectively. These large-scale differences may imply that there is an absent large-scale effect in the models, although we reserve further analysis to future work with a larger-volume simulation.